OpenStudy (anonymous):

~ Mark is twice as old as his sister Jade. Jade is (x) years old . Their sister Marika is three years older than Mark . a) Write mark's age , in terms of x . b) Write marika's age in terms of x . c) Find in terms of x , the sum of their ages .

OpenStudy (anonymous):

mark's age =2x marika's age = 2x+3 sum of ages= x+2x+2x+3 = 5x+3

OpenStudy (anonymous):

thankyou very much :D